# Semantic Component Analysis: Discovering Patterns in Short Texts Beyond Topics

### v1

@@ 8.2 SCA as a Transformation
$\mathbf{x}^{\prime}_{0}=\mathbf{x},\\ \mathbf{x}^{\prime}_{i+1}=\mathbf{x}^{\prime}_{i}-\mu\mathbb{1}_{\alpha_{i,i}>\alpha}\langle\mathbf{x}_{i},\mathbf{v}_{i}\rangle\mathbf{v}_{i}$ (3)
@@
$a_{j}=\mu\mathbb{1}_{\alpha_{i,i}>\alpha}\langle\mathbf{x}^{\prime}_{j},\mathbf{v}_{j}\rangle$ (4)
@@ 8.3 Stopping Criteria
9 Extended Results
9.1 Implementation and Preprocessing
The datasets we use for evaluating our method have been obtained from different sources. The Tweets made by Donald Trump are available in the Trump Twitter Archive Brown (2021), from where we download the dataset directly as .csv which resulted in 60,053 samples as of 10th of December 2023. The Chinese News Tweets have been scraped from a list of Chinese news outlets, which we compiled by two criteria: (1) The majority of the Tweets uses simplified or traditional Chinese characters and (2) the news outlet is based outside of the borders of the People’s Republic of China. The list is far from complete, but we stopped adding new accounts when a threshold of 1M Tweets was reached. The accounts included are: “voachinese", “chinesewsj", “abcchinese", “reuterschina", “dw_chinese", “rfi_tradcn", “asiafinance", “initiumnews", “cdtchinese", “kbschinese", “rijingzhongwen", “weiquanwang", “zaobaosg", “rfa_chinese", “rfi_cn", “rcizhongwen", “kbschines", and “chosunchinese".
@@ 9.2 Evaluation Metrics
  • Noise rate: The fraction of documents that are not assigned to any cluster in the first iteration (1st) and over all iterations.
@@
Figure 4: Run statistics for the three datasets. (1st) denotes the statistic calculated only over the clusters/components of the first iteration of SCA, which is equivalent to a BERTopic run with the same hyperparameters.
@@
Figure 5: Grid run results for different hyperparameters. All configurations were run with $\theta=1.0$ to avoid occluding “bad" results by merging many components.
9.3 Trump Dataset: Components
@@
“A very good way to pave your own way to success is simply to work hard and to be diligent” – Think Like a Champion
@@
The New York Times has apologized for the terrible Anti-Semitic Cartoon, but they haven’t apologized to me for this or all of the Fake and Corrupt news they print on a daily basis. They have reached the lowest level of “journalism,” and certainly a low point in @nytimes history! 0.22
@@
"""It’s more important to be smart than tough. I know businessmen who are brutally tough, but they’re not smart.” – Think Like A Billionaire" 0.05
@@
9.4 Chinese News Outlet Tweets: Components
@@
9.5 Hausa Dataset: Components
@@
9.6 Performance of Transformed Embeddings on MTEB
@@
As a linear decomposition, SCA is a transformation of the embeddings into an explainable space, where each dimension corresponds to the score along one semantic component as defined by formula (2) (see Appendix 8.2 for details). We use this interpretation of the component scores to evaluate the topic distribution in terms of its performance as a representation of samples in downstream tasks defined in the MTEB benchmark (Muennighoff et al., 2023) and compare against the vanilla embeddings as well as a PCA transformation of the same dimension.

### v2

@@ A.2 SCA as a Transformation
$\mathbf{x}^{\prime}_{0}=\mathbf{x},\\ \mathbf{x}^{\prime}_{i+1}=\mathbf{x}^{\prime}_{i}-\mu\mathbbm{1}_{\alpha_{i,i}>\alpha}\langle\mathbf{x}_{i},\mathbf{v}_{i}\rangle\mathbf{v}_{i}$ (3)
@@
$a_{j}=\mu\mathbbm{1}_{\alpha_{i,i}>\alpha}\langle\mathbf{x}^{\prime}_{j},\mathbf{v}_{j}\rangle$ (4)
@@ A.3 Stopping Criteria
Appendix B Data & Evaluation Appendix
B.1 Implementation and Preprocessing
The datasets we use for evaluating our method have been obtained from different sources. The Tweets made by Donald Trump are available in the Trump Twitter Archive Brown (2021), from where we download the dataset directly as .csv which resulted in 60,053 samples as of 10th of December 2023. The Chinese News Tweets have been scraped from a list of Chinese news outlets, which we compiled by two criteria: (1) The majority of the Tweets uses simplified or traditional Chinese characters and (2) the news outlet is based outside of the borders of the People’s Republic of China. The list is far from complete, but we stopped adding new accounts when a threshold of 1M Tweets was reached. The accounts included are: "voachinese", "chinesewsj", "abcchinese", "reuterschina", "dw_chinese", "rfi_tradcn", "asiafinance", "initiumnews", "cdtchinese", "kbschinese", "rijingzhongwen", "weiquanwang", "zaobaosg", "rfa_chinese", "rfi_cn", "rcizhongwen", "kbschines", and "chosunchinese".
@@ B.2 Evaluation Metrics
  • Noise rate: The fra¥ction of documents that are not assigned to any cluster in the first iteration (1st) and over all iterations.
@@ C.1 Run Statistics and Grid Run
Figure 3: Run statistics for the three datasets. (1st) denotes the statistic calculated only over the clusters/components of the first iteration of SCA, which is equivalent to a BERTopic run with the same hyperparameters. Table 1 provides the same values, but we keep this table here for reference mirroring the layout of our original submission.
@@
Figure 4: Grid run results for different hyperparameters. All configurations were run with $\theta=1.0$ to avoid occluding "bad" results by merging many components.
C.2 Trump Dataset: Components
@@
"A very good way to pave your own way to success is simply to work hard and to be diligent" - Think Like a Champion
@@
The New York Times has apologized for the terrible Anti-Semitic Cartoon, but they haven’t apologized to me for this or all of the Fake and Corrupt news they print on a daily basis. They have reached the lowest level of "journalism," and certainly a low point in @nytimes history! 0.22
@@
"""It’s more important to be smart than tough. I know businessmen who are brutally tough, but they’re not smart." - Think Like A Billionaire" 0.05
@@
C.3 Chinese News Outlet Tweets: Components
@@
C.4 Hausa Dataset: Components
@@
C.5 Performance of Transformed Embeddings on MTEB
@@
As a linear decomposition, SCA is a transformation of the embeddings into an explainable space, where each dimension corresponds to the score along one semantic component as defined by formula (2) (see Appendix A.2 for details). We use this interpretation of the component scores to evaluate the topic distribution in terms of its performance as a representation of samples in downstream tasks defined in the MTEB benchmark (Muennighoff et al., 2023) and compare against the vanilla embeddings as well as a PCA transformation of the same dimension.